Stacey is selling tickets to the school play. The tickets are $6 for adults and $5 for children. She sells twice as many adult tickets as children's tickets and brings in a total of $391. How many of each kind of ticket did she sell?

Answer by ikleyn(52834)

.

Let x be the number of children's tickets; then the number of adult tickets is 2x.


The revenue equation is then


    5x + 6*(2x) = 391.


Simplify and solve


    5x + 12x   = 391

    17x        = 391

      x        = 391/17 = 23.


ANSWER.  23 children's tickets and 2*23 = 46 adult tickets.


CHECK.   5*23 + 6*46 = 391 dollars.   ! Correct !

Solved.